OpenStudy (anonymous):

The circumference of a circle is 8 inches. Find its radius and area the radius of the circle is ? inches

OpenStudy (anonymous):

Radius = 8 inches Circumference of the Circle = 2*(3.14)*8 inches = 16*(3.14) inches = 50.24 inches So... it's 50.24

OpenStudy (anonymous):

Click "Best answer" if my answer helped. :)

OpenStudy (anonymous):

@Sean_Le_Van the circumference is 8 inches, not the radius.

OpenStudy (anonymous):

\[C=2\pi r\]\[8=2 \pi r\]\[r=\frac{8}{2\pi}\]\[r \approx 1.27 in \]

OpenStudy (anonymous):

The area is:\[A=\pi r^2\]\[A=\pi (\frac{8}{2 \pi})^2\]\[A=\frac{64}{4\pi}\]\[A \approx 5.09 in^2\]